# LITP-25-17

### v1

@@ 1 Introduction
Scattering amplitudes lie at the core of both theoretical and experimental high-energy physics. They connect directly to physical observables accessible in experiments and at the same time encode the principles of quantum mechanics and relativity in a remarkably compact form. For this reason, amplitudes play a central role in the S-matrix bootstrap program, through which causality and unitarity can define the space of UV-completable effective field theories (EFTs) through positivity constraints, see e.g. [Adams:2006sv, Camanho:2014apa, Pham:1985cr, Pennington:1994kc, Ananthanarayan:1994hf, Distler:2006if, Vecchi:2007na, Manohar:2008tc, Bellazzini:2014waa, Bellazzini:2015cra, Bellazzini:2016xrt, Cheung:2016yqr, Bonifacio:2016wcb, Hinterbichler:2017qcl, deRham:2017avq, deRham:2017zjm, Bellazzini:2017fep, Bellazzini:2018paj, Bellazzini:2019bzh, Bellazzini:2019xts, Alberte:2020bdz, Englert:2019zmt, Melville:2019tdc, Bellazzini:2020cot, Arkani-Hamed:2020blm, Caron-Huot:2020cmc, Tolley:2020gtv, Bellazzini:2021oaj, Bern:2021ppb, Chiang:2021ziz, Caron-Huot:2021rmr, Chala:2021wpj, Davighi:2021osh, Arkani-Hamed:2021ajd, Bellazzini:2021shn, Serra:2022pzl, Caron-Huot:2022jli, Bellazzini:2022wzv, Caron-Huot:2022ugt, Guerrieri:2022sod, Henriksson:2022oeu, Haring:2022sdp, Creminelli:2022onn, Bellazzini:2023nqj, Karateev:2023mrb, Haring:2023zwu, McCullough:2023szd, Albert:2023jtd, Ma:2023vgc, Berman:2023jys, EliasMiro:2023fqi, Albert:2023seb, Creminelli:2023kze, Berman:2024kdh, Albert:2024yap, Berman:2024eid, Berman:2024owc, Creminelli:2024lhd, Dong:2024omo, Bertucci:2024qzt, Cremonini:2024lxn, Remmen:2024hry, Haring:2024wyz, Tokareva:2025rta, Cheung:2025krg, deRham:2025vaq, Peng:2025klv, Chang:2025cxc, Hui:2025aja, Correia:2025uvc, Ye:2025zhs, Berman:2025owb, Bellazzini:2025shd, Bonnefoy:2025uzf, Dong:2025dpy, Creminelli:2025rxj, Huang:2025icl].
@@
This perspective is challenged in four spacetime dimensions by the existence of long-range forces mediated by photons and gravitons, whose infrared (IR) divergences trivialize scattering amplitudes, driving them to zero. [^1] An equivalent viewpoint on IR divergences has emerged in the context of memory effect [Satishchandran:2019pyc, Prabhu:2022zcr] and asymptotic symmetries—large gauge transformations in QED and BMS symmetries in gravity—where amplitudes vanish because they violate selection rules associated with asymptotic charges [Gabai:2016kuf, Kapec:2017tkm, Strominger:2017zoo]. IR divergences are tamed in inclusive observables, as guaranteed by the Bloch–Nordsieck and KLN theorems [Bloch:1937pw, Kinoshita:1962ur, Lee:1964is], as well as in certain IR-finite quantities such as energy–energy correlators [Moult:2025nhu]. Yet these observables have not been successfully incorporated into the bootstrap program, leading to an apparent breakdown of amplitude constraints as soon as long-range mediators are present, no matter how weakly they couple. Although this discontinuity seems paradoxical, it is natural to wonder whether such extreme UV/IR mixing, perhaps reminiscent of the swampland program [Vafa:2005ui, ArkaniHamed:2006dz, Palti:2019pca], might be unavoidable.
@@
In this paper, we address this specific question, showing that one can construct bounds that smoothly interpolate between finite and vanishing couplings in electromagnetism and gravity. The key observation is that long-range interactions require initial and final scattering states to be specified, at least partially, in terms of properties of an experimental apparatus. Unlike in the case of short-range interactions or IR-safe observables, the apparatus cannot be fully decoupled from the scattering process under study. This is reflected, for example, in the infinitely many viable choices of asymptotic Hamiltonians and dressings in the Faddeev–Kulish approach to IR-finite amplitudes [Kulish:1970ut, Kapec:2017tkm, Hannesdottir:2019opa].
@@
We focus on experimental apparatuses of finite size $\mathcal{E}^{-1}$ which, in line with the traditional definition of amplitudes, we take exponentially larger than any intrinsic distance-scale $1/M$ of the system under study: $\log(M/\mathcal{E})\gg 1$. The ability of the apparatus to detect low-energy (soft) particles is limited by its finite inverse size. In these conditions, the most important effects are captured by the scaling limit,
@@
with $G=1/8\pi m^{2}_{\mathrm{Pl}}$ the Newton constant, $m_{\mathrm{Pl}}$ the reduced Planck mass, and $\alpha=e^{2}/4\pi$ the fine structure constant. Crucially, the theory effectively depends on two couplings, $\alpha$ and $\alpha_{\mathcal{E}}$ ($GM^{2}$ and $G_{\mathcal{E}}$) —in addition to any other short–range interactions— and these couplings play qualitatively and quantitatively separate roles. The scaling limit isolates the universal long–distance effects common to all sufficiently large detectors, in terms of a single parameter, $\alpha_{\mathcal{E}}$ ($G_{\mathcal{E}}$). Finite corrections from the couplings $\alpha$ and $G$, as well as finer dependence on $\mathcal{E}$ and other details of the experimental apparatus, provide subleading contributions to observables, entering only at next–to–leading order in the scaling (1.1).  [^2] Perhaps the famous quote by V.I. Arnold “Mathematics is the part of physics where experiments are cheap” [Arnold:98] helps expressing our perspective on the physical interpretation of the mathematical limit (1.1).
We then define stripped amplitudes -$\mathcal{M}_{\mathcal{E}}$ as a formal amplitude in the finite-size experiment, by factoring out the analytic extension of Weinberg IR–divergent exponentials [Weinberg:1965nx]. We prove that the stripped amplitudes are IR-finite, analytic, crossing symmetric, Lorentz invariant, and Regge–behaved. We moreover show that they are unitary in the scaling limit (1.1), and that their exclusive cross sections reproduce the IR–finite inclusive ones controlled by the energy–resolution scale $\mathcal{E}$. The universality of this leading behavior also ensures that in the same scaling limit they match Faddeev–Kulish amplitudes.
These features allow dispersive methods to be extended to theories with long-range interactions and derive positivity bounds on EFT Wilson coefficients from dispersive sum rules for $2\to 2$ stripped amplitudes. As explicit examples, we establish IR–finite positivity bounds in the EFT of pions coupled to electromagnetism or gravity. Notably, these are also free of the issues that arise due to the photon/graviton $1/t$-pole in four dimensions [Caron-Huot:2021rmr, Henriksson:2022oeu]. Moreover, they resolve the puzzle of non-decoupling gravitational negativity [Chang:2025cxc], showing that only a finite, decoupling contribution $\sim GM^{2}\log M/\mathcal{E}$ remains, while robust inequalities hold in $D=4$.
@@ 2 Stripped and Hard Amplitudes
We next introduce a scale $\mathcal{E}>0$ and call photons and gravitons “soft” if they have energy below $\mathcal{E}$ or “hard” otherwise. As the IR regulator $\epsilon$ is removed, the probability for real soft emission increases, and the survival probability of exclusive amplitudes among states with only hard particles, $|\boldsymbol{k}|>\mathcal{E}$, decays exponentially. In practice, this suppression arises from loops of virtual soft particles with $|\boldsymbol{k}|<\mathcal{E}$, whose singular contributions can be computed, since they exponentiate. In QED, the exclusive (regulated) amplitude $\mathcal{M}^{\epsilon}$ decays exponentially with the overall factor obtained by adapting the original result by Weinberg [Weinberg:1965nx],
$\mathcal{W}^{\mathrm{QED}}=\mathrm{exp}\left\{\frac{\alpha}{8\pi}\frac{\left(\mathcal{E}/\mu\right)^{2\epsilon}}{\epsilon}\left(\sum_{i,j}\eta_{i}\eta_{j}q_{i}q_{j}\frac{1}{\beta_{ij}}\log\frac{1+\beta_{ij}}{1-\beta_{ij}}-i\sum_{\begin{subarray}{c}i\neq j\\ \eta_{i}\eta_{j}=1\end{subarray}}q_{i}q_{j}\frac{2\pi}{\beta_{ij}}+O(\epsilon)\right)\right\}\,$, (2.2)
where $\eta_{i}=\pm 1$ for in (-) and out (++) going states, $q_{i}$ are the in/out electric charges, [^4] While this expression is for in/out notation, in Section 3 we will compute its analytic continutation, which will enable us to work with all-incoming notation. and $\beta_{ij}\equiv[1-(m_{i}m_{j}/k_{i}\!\cdot\!k_{j})^{2}]^{1/2}$ the relative velocity between particles $i$ and $j$; $\mu$ is the dimensional-regularization scale, which eventually drops out of all physical expressions and the $O(\epsilon)$ term is $(\mathcal{E}/\mu)$-independent. As advertised, $\lim_{\epsilon\to 0^{+}}\mathcal{W}^{\mathrm{QED}}=0$, while if we first take $\alpha\to 0$ the amplitude is finite as $\epsilon\to 0$.
Gravitational effects lead to a similar exponential suppression [Weinberg:1965nx],
$\mathcal{W}^{\mathrm{GR}}\!=\!\mathrm{exp}\left\{\!-\frac{G}{8\pi}\frac{\left(\mathcal{E}/\mu\right)^{2\epsilon}}{\epsilon}\!\!\left(\sum_{i,j}\eta_{i}\eta_{j}m_{i}m_{j}\frac{1+\beta^{2}_{ij}}{\beta_{ij}\sqrt{1-\beta^{2}_{ij}}}\log\frac{1+\beta_{ij}}{1-\beta_{ij}}-i\!\!\sum_{\begin{subarray}{c}i\neq j\\ \eta_{i}\eta_{j}=1\end{subarray}}m_{i}m_{j}\frac{2\pi(1+\beta^{2}_{ij})}{\beta_{ij}\sqrt{1-\beta^{2}_{ij}}}+O(\epsilon)\!\!\right)\!\!\right\}\,$, (2.3)
@@
Motivated by the universal factorization of IR–divergent soft contributions, we isolate the contribution from processes with hard kinematics, $|s_{ij}|>\mathcal{E}^{2}$, by first defining the ratio obtained by stripping off $\mathcal{W}$ (i.e. $\mathcal{W}^{\mathrm{QED}}$, $\mathcal{W}^{\mathrm{GR}}$ or both) from the full amplitude:
@@
Since, by construction, $\mathcal{M}_{\mathcal{E}}^{\epsilon}$ is IR finite, the stripped amplitude $\mathcal{M}_{\mathcal{E}}$ depends only on the physical scale $\mathcal{E}$ and kinematics. In perturbation theory, $\mathcal{M}_{\mathcal{E}}$ receives contributions from loops of virtual photons/gravitons that are either hard $|\boldsymbol{k}|>\mathcal{E}$, or soft with $|\boldsymbol{k}|<\mathcal{E}$ but finite (since the singular part has been canceled by $\mathcal{W}$). We write these as
@@
[…]
@@ 3.1 Analytic soft exponentials
Therefore, as $s_{ij}\to\infty$, the exponent goes to a constant, meaning $\mathcal{W}^{\mathrm{QED}}_{ij}/s_{ij}\to 0$ in this limit. We can then use once-subtracted dispersion relations to analytically continue the Weinberg exponentials, [^7] An alternative route is going through a worldline representation, see [Lippstreu:2025jit].
$\mathcal{W}^{\mathrm{QED}}_{ij}(s_{ij})=\mathcal{W}^{\mathrm{QED}}_{ij}(s_{0})+\int_{(m_{1}+m_{2})^{2}}^{\infty}\frac{d\bar{s}}{2\pi i}\frac{s_{ij}-s_{0}}{(\bar{s}-s_{0})(\bar{s}-s_{ij})}\mathrm{Disc}\,\mathcal{W}^{\mathrm{QED}}_{ij}(\bar{s})\,$, (3.3)
@@
which are, by construction, analytic functions with no branch point at the pseudothreshold $(m_{i}-m_{j})^{2}$, nor at any unphysical location. For real kinematics, their product matches (2.2). For illustration, with $m_{i}^{2}=m_{j}^{2}=m\ll|s_{ij}|$, these factors simplify to,
$\mathcal{W}^{\mathrm{QED}}_{ij}(s_{ij})\to\mathrm{exp}\left[\frac{\alpha q_{i}q_{j}}{4\pi}\frac{\left(\mathcal{E}/\mu\right)^{2\epsilon}}{\epsilon}\log\left(-\frac{s_{ij}}{m^{2}}\right)\right]\,,\quad\quad(m\ll|s_{ij}|\,,i\neq j)\,.$ (3.5)

### v2

@@ 1 Introduction
Scattering amplitudes lie at the core of both theoretical and experimental high-energy physics. They connect directly to physical observables accessible in experiments and at the same time encode the principles of quantum mechanics and relativity in a remarkably compact form. For this reason, amplitudes play a central role in the S-matrix bootstrap program, through which causality and unitarity can define the space of UV-completable effective field theories (EFTs) through positivity constraints, see e.g. [1, 44, 127, 126, 9, 69, 139, 114, 24, 18, 30, 54, 42, 100, 66, 67, 27, 29, 28, 23, 8, 78, 117, 19, 11, 50, 137, 26, 37, 56, 49, 52, 65, 12, 20, 131, 47, 22, 46, 87, 98, 92, 62, 21, 105, 94, 116, 7, 113, 34, 75, 4, 60, 36, 5, 33, 35, 61, 70, 38, 64, 129, 95, 51, 135, 55, 68, 125, 53, 103, 58, 143, 32, 25, 43, 71, 63, 102].
@@
This perspective is challenged in four spacetime dimensions by the existence of long-range forces mediated by photons and gravitons, whose infrared (IR) divergences trivialize scattering amplitudes, driving them to zero. [^1] An equivalent viewpoint on IR divergences has emerged in the context of memory effect [130, 128] and asymptotic symmetries—large gauge transformations in QED and BMS symmetries in gravity—where amplitudes vanish because they violate selection rules associated with asymptotic charges [83, 104, 134]. IR divergences are tamed in inclusive observables, as guaranteed by the Bloch–Nordsieck and KLN theorems [41, 107, 110], as well as in certain IR-finite quantities such as energy–energy correlators [118]. Yet these observables have not been successfully incorporated into the bootstrap program, leading to an apparent breakdown of amplitude constraints as soon as long-range mediators are present, no matter how weakly they couple. Although this discontinuity seems paradoxical, it is natural to wonder whether such extreme UV/IR mixing, perhaps reminiscent of the swampland program [138, 13, 121], might be unavoidable.
@@
In this paper, we address this specific question, showing that one can construct bounds that smoothly interpolate between finite and vanishing couplings in electromagnetism and gravity. The key observation is that long-range interactions require initial and final scattering states to be specified, at least partially, in terms of properties of an experimental apparatus. Unlike in the case of short-range interactions or IR-safe observables, the apparatus cannot be fully decoupled from the scattering process under study. This is reflected, for example, in the infinitely many viable choices of asymptotic Hamiltonians and dressings in the Faddeev–Kulish approach to IR-finite amplitudes [109, 104, 91].
@@
We focus on experimental apparatuses of finite proper size $\mathcal{E}^{-1}$ which, in line with the traditional definition of amplitudes, we take exponentially larger than any intrinsic distance-scale $1/M$ of the system under study: $\log(M/\mathcal{E})\gg 1$. The ability of the apparatus to detect low-energy (soft) particles is limited by its finite inverse size. In these conditions, the most important effects are captured by the scaling limit,
@@
with $G=1/8\pi m^{2}_{\mathrm{Pl}}$ the Newton constant, $m_{\mathrm{Pl}}$ the reduced Planck mass, and $\alpha=e^{2}/4\pi$ the fine structure constant. Crucially, the theory effectively depends on two couplings, $\alpha$ and $\alpha_{\mathcal{E}}$ ($GM^{2}$ and $G_{\mathcal{E}}$) —in addition to any other short–range interactions— and these couplings play qualitatively and quantitatively separate roles. The scaling limit isolates the universal long–distance effects common to all sufficiently large detectors, in terms of a single parameter, $\alpha_{\mathcal{E}}$ ($G_{\mathcal{E}}$). Finite corrections from the couplings $\alpha$ and $G$, as well as finer dependence on $\mathcal{E}$ and other details of the experimental apparatus, provide subleading contributions to observables, entering only at next–to–leading order in the scaling (1.1).  [^2] Perhaps the famous quote by V.I. Arnold “Mathematics is the part of physics where experiments are cheap” [14] helps expressing our perspective on the physical interpretation of the mathematical limit (1.1).
We then define stripped amplitudes $\mathcal{M}_{\mathcal{E}}$ as a formal amplitude in the finite-size experiment, by factoring out the analytic extension of Weinberg IR–divergent exponentials [141]. We prove that the stripped amplitudes are IR-finite, analytic, crossing symmetric, Lorentz invariant, and Regge–behaved. We moreover show that they are unitary in the scaling limit (1.1), and that their exclusive cross sections reproduce the IR–finite inclusive ones controlled by the energy–resolution scale $\mathcal{E}$. The universality of this leading behavior also ensures that in the same scaling limit they match Faddeev–Kulish amplitudes.
These features allow dispersive methods to be extended to theories with long-range interactions and derive positivity bounds on EFT Wilson coefficients from dispersive sum rules for $2\to 2$ stripped amplitudes. As explicit examples, we establish IR–finite positivity bounds in the EFT of pions coupled to electromagnetism or gravity. Notably, these are also free of the issues that arise due to the photon/graviton $1/t$-pole in four dimensions [49, 98]. Moreover, they resolve the puzzle of non-decoupling gravitational negativity [53], showing that only a finite, decoupling contribution $\sim GM^{2}\log M/\mathcal{E}$ remains, while robust inequalities hold in $D=4$.
As a simple illustration, consider the EFT of a single pion with $2\to 2$ scattering amplitude $\mathcal{M}=\hat{c}_{2,0}(s^{2}+t^{2}+u^{2})/2+\ldots$. In the absence of long–range forces, positivity implies the familiar bound $\hat{c}_{2,0}\geq 0$. In the presence of gravity, the bound takes the following form
$\mathcal{A}_{0}(\hat{c}_{n,k},G_{\mathcal{E}})+O(G)\geq 0\,$, (1.2)
with $\mathcal{A}_{0}$ a known function of Wilson coefficients and of $G_{\mathcal{E}}$. In other words, we obtain finite and calculable corrections in $D=4$ to the usual positivity constraints, unlike previous analyses. For instance, a representative (and not optimized) bound obtained in our analysis reads $\hat{c}_{2,0}M^{4}+65.7\!\left(8\pi G_{\mathcal{E}}+O(G_{\mathcal{E}}^{2})\right)+O(G)>0$, with $G_{\mathcal{E}}$ encoding the universal dominant long–range effect of gravity, which is systematically calculable (see (7.5) and (7.6) for the explicit $O(G_{\mathcal{E}}^{2})$ terms). The smallness of $G/G_{\mathcal{E}}\ll 1$ ensures parametric control over the $O(G)$ corrections.
@@ 2 Stripped and Hard Amplitudes
We next introduce a Lorentz invariant scale $\mathcal{E}>0$, which can correspond to the inverse proper size of an experiment, or its (proper) energy resolution, or even just a reference mass of a particle. To isolate the most important IR effects, we momentarily pick the rest frame of the experiment and label photons and gravitons “soft” if they have energy below $\mathcal{E}$ or “hard” otherwise. [^4] This intermediate step is effectively defining the Lorentz breaking hard amplitude $\mathcal{M}^{\mathrm{hard}}_{\mathcal{E}}$ that is introduced in (2.5), from which we then abstract to the Lorentz covariant definition of stripped amplitudes in (2.4) or (3.8). As the IR regulator $\epsilon$ is removed, the probability for real soft emission increases, and the survival probability of exclusive amplitudes among states with only hard particles, $|\boldsymbol{k}|>\mathcal{E}$, decays exponentially. In practice, this suppression arises from loops of virtual soft particles with $|\boldsymbol{k}|<\mathcal{E}$, whose singular contributions can be computed, since they exponentiate. In QED, the exclusive (regulated) amplitude $\mathcal{M}^{\epsilon}$ decays exponentially with the overall factor obtained by adapting the original result by Weinberg [141],
$\mathcal{W}^{\mathrm{QED}}\equiv\mathrm{exp}\left\{\frac{\alpha}{8\pi}\frac{\left(\mathcal{E}/\mu\right)^{2\epsilon}}{\epsilon}\left(\sum_{i,j}\eta_{i}\eta_{j}q_{i}q_{j}\frac{1}{\beta_{ij}}\log\frac{1+\beta_{ij}}{1-\beta_{ij}}-i\sum_{\begin{subarray}{c}i\neq j\\ \eta_{i}\eta_{j}=1\end{subarray}}q_{i}q_{j}\frac{2\pi}{\beta_{ij}}\right)\right\}\,$, (2.2)
where $\eta_{i}=\pm 1$ for in (-) and out (++) going states, $q_{i}$ are the in/out electric charges, [^5] While this expression is for in/out notation, in Section 3 we will compute its analytic continuation, which will enable us to work with all-incoming notation. and $\beta_{ij}\equiv[1-(m_{i}m_{j}/k_{i}\!\cdot\!k_{j})^{2}]^{1/2}$ the relative velocity between particles $i$ and $j$; $\mu$ is the dimensional-regularization scale, which eventually drops out of all physical expressions. Importantly, in the definition of the overall factor $\mathcal{W}^{\mathrm{QED}}$ we singled out the universal Lorentz invariant contribution, despite the intermediate steps involved a frame–dependent labeling for photons. As advertised, $\lim_{\epsilon\to 0^{+}}\mathcal{W}^{\mathrm{QED}}=0$, while if we first take $\alpha\to 0$ the amplitude is finite as $\epsilon\to 0$.
Gravitational effects lead to a similar exponential suppression [141],
$\mathcal{W}^{\mathrm{GR}}\!\equiv\!\mathrm{exp}\left\{\!-\frac{G}{8\pi}\frac{\left(\mathcal{E}/\mu\right)^{2\epsilon}}{\epsilon}\!\!\left(\sum_{i,j}\eta_{i}\eta_{j}m_{i}m_{j}\frac{1+\beta^{2}_{ij}}{\beta_{ij}\sqrt{1-\beta^{2}_{ij}}}\log\frac{1+\beta_{ij}}{1-\beta_{ij}}-i\!\!\sum_{\begin{subarray}{c}i\neq j\\ \eta_{i}\eta_{j}=1\end{subarray}}m_{i}m_{j}\frac{2\pi(1+\beta^{2}_{ij})}{\beta_{ij}\sqrt{1-\beta^{2}_{ij}}}\!\!\right)\!\!\right\},$ (2.3)
@@
Motivated by the universal factorization of IR–divergent soft contributions, we define the following ratio obtained by stripping off $\mathcal{W}$ (i.e. $\mathcal{W}^{\mathrm{QED}}$, $\mathcal{W}^{\mathrm{GR}}$ or both) from the full amplitude:
@@
Since, by construction, $\mathcal{M}_{\mathcal{E}}^{\epsilon}$ is IR finite, the stripped amplitude $\mathcal{M}_{\mathcal{E}}$ depends only on the physical scale $\mathcal{E}$ and kinematics.
[…]
@@
The hard amplitude $\mathcal{M}^{\mathrm{hard}}_{\mathcal{E}}$ in (2.5) can be independently defined by a path integral in which the original Lagrangian $\mathcal{L}(A^{\mu},h^{\mu\nu},\phi_{i})$ is replaced by $\mathcal{L}(A^{\mu}_{\mathcal{E}},h^{\mu\nu}_{\mathcal{E}},\phi_{i})$, where the gauge and gravitational fields $A^{\mu}_{\mathcal{E}}$ and $h^{\mu\nu}_{\mathcal{E}}$ contain no Fourier modes with $|\boldsymbol{k}|<\mathcal{E}$. In particular, all photon and graviton propagators used to perturbatively build $\mathcal{M}^{\mathrm{hard}}_{\mathcal{E}}$ come with a step function $\theta(|\boldsymbol{k}|-\mathcal{E})$, and the associated interaction doesn’t include vertices with soft photons/gravitons. On the other hand, $\delta_{\mathrm{sub.~soft}}$ contains the contributions from soft photon or graviton loops which are not captured by the overall divergent exponential $\mathcal{W}$ because they are IR-finite and subleading in the $\mathcal{E}\to 0$ limit. Since the IR regulated amplitude $\mathcal{M}^{\epsilon}$ is independent of $\mathcal{E}$, the stripped amplitude $\mathcal{M}_{\mathcal{E}}$ depends only on $\log\mathcal{E}$ via $\mathcal{W}$, whereas $\mathcal{M}_{\mathcal{E}}^{\mathrm{hard}}$ and $\delta_{\mathrm{sub.~soft}}$ may separately depend also on powers of $\mathcal{E}/E\ll 1$, which cancel out in their sum. [^7] Ref. [141] implies $\mathcal{M}^{\epsilon}=\mathcal{M}^{\mathrm{hard}}_{\mathcal{E}}\mathcal{W}+\text{subleading soft}$. However, all evidence from exactly solvable models [41, 39] as well as from conservation of asymptotic charges [134] and path integral resummations [142], indicates that the exponential $\mathcal{W}$ multiplies the full amplitude as a global suppression factor, rather than acting only as a prefactor to the hard part. We adopt this physically safe assumption throughout this work.
@@ 3.1 Analytic soft exponentials
Therefore, as $s_{ij}\to\infty$, the $\mathcal{W}^{\mathrm{QED}}_{ij}/s_{ij}\to 0$. We can then use once-subtracted dispersion relations to analytically continue the Weinberg exponentials, [^8] An alternative route is going through a worldline representation, see [112].
$\mathcal{W}^{\mathrm{QED}}_{ij}(s_{ij})=\mathcal{W}^{\mathrm{QED}}_{ij}(s_{0})+\int_{(m_{i}+m_{j})^{2}}^{\infty}\frac{d\bar{s}}{2\pi i}\frac{s_{ij}-s_{0}}{(\bar{s}-s_{0})(\bar{s}-s_{ij})}\mathrm{Disc}\,\mathcal{W}^{\mathrm{QED}}_{ij}(\bar{s})\,$, (3.3)
@@
which are, by construction, analytic functions with no branch point at the pseudothreshold $(m_{i}-m_{j})^{2}$, nor at any unphysical location. For real kinematics, their product matches (2.2). For illustration, with $m_{i}^{2}=m_{j}^{2}=m^{2}\ll|s_{ij}|$, these factors simplify to,
$\mathcal{W}^{\mathrm{QED}}_{ij}(s_{ij})\to\mathrm{exp}\left[\frac{\alpha q_{i}q_{j}}{4\pi}\frac{\left(\mathcal{E}/\mu\right)^{2\epsilon}}{\epsilon}\log\left(-\frac{s_{ij}}{m^{2}}\right)\right]\,,\quad\quad(m^{2}\ll|s_{ij}|\,,i\neq j)\,.$ (3.5)